OpenStudy (anonymous):

The vertex form of the equation of a parabola is y = 3(x - 4)2 - 22. What is the standard form of the equation? Will medal!

OpenStudy (anonymous):

A. y = 3x2 - 24x + 26 B. y = 9x2 - 26 C. y = 3x2 - 24x + 4 D. y = 3x2 - 18 ^ these are the answer choices. :)

OpenStudy (ddcamp):

The standard form is \[y = ax^2 + bx + c\] To figure out which of the answers is correct, let's start by expanding the (x-4)² term.

OpenStudy (ddcamp):

What's \[3*(x-4)*(x-4)?\]

OpenStudy (anonymous):

I know that x-4 expanded is 8x. 8x times 3 = 24x :3

OpenStudy (ddcamp):

How did you get 8x from (x-4)*(x-4)?

OpenStudy (anonymous):

I multiplied the two 4x's. It's a bit of a shortcut, but it helps. :p

OpenStudy (ddcamp):

I don't think that shortcut helps here. Let's try the good ol' FOIL method: